OpenStudy (anonymous):

how do you solve 5/8 y=1/2

OpenStudy (anonymous):

\[5y=(1/2)*8\] \[5y = 4\] \[y=4/5\]

OpenStudy (anonymous):

5\8y=1\2 =)y=1\2*8\5 =)y=4\5

jhonyy9 (jhonyy9):

check the common denominator what is 8y - so and after you will get 5=4y from what y=5/4

jhonyy9 (jhonyy9):

or if you want to write (5y)/8 = 1/2 so than you get 5y=4 and y= 4/5